Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2024 logbook, Contact: F.Wolff-Fabris, A.Galler, H.Sinn  Not logged in ELOG logo
Message ID: 133     Entry time: 24 Jul 2024, 15:22
Author: Giuseppe Mercurio 
Group: SCS 
Subject: Shift summary 

Shift summary (Natalia, Giuseppe)

Achievements:

  • M1, M2 pitches adjusted
  • Grating AX changed value during shutdown, encoder offsets have been recalibrated
  • Set up SCS, 9mrad, HEPM, grating configuration
  • Set up SCS, 9 mrad, HEPM, blank configuration

Problems:

  • at 10am the mono translations gone in error. It was problem with fuse in ELMO crate in tunnel. Repaired (ZZ). Machine was log to set up, till end of shift, we lost ~5 hours from the shift.
ELOG V3.1.4-7c3fd00